# Kronecker factorization theorems for the exceptional Malcev algebra

### v2

@@ Abstract
We prove that a Malcev algebra $\mathcal{M}$ containing the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}A)\oplus\mathcal{N}$, where $A$ is a commutative associative algebra. Also, we prove that a Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ whose even part $\mathcal{M}_{0}$ contains $\mathbb{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}\mathcal{A})\oplus\mathcal{N^{\prime}}$, where $\mathcal{A}$ is a supercommutative associative superalgebra.
@@ 1 Introduction
In the case of superalgebras, M. López-Díaz and I. Shestakov [10, 9] studied the representations of simple alternative and exceptional Jordan superalgebras in characteristic $3$ and through these representations, they obtained some analogues of the Kronecker Factorization Theorem for these superalgebras. Also, the author [11] obtained some analogues of the Kronecker Factorization Theorem for central simple alternative superalgebras $M_{(1\mid 1)}(F)$, $\mathbb{O}(4|4)$ and $\mathbb{O}[u]$, where in particular the Kronecker Factorization Theorem for $M_{(1\mid 1)}(F)$ answers the analogue for superalgebras of the Jacobson’s problem [6]. We state that Jacobson’s problem [6] was recently solved by the author and I. Shestakov [12] in the split case. Similarly, C. Martinez and E. Zelmanov [13] obtained a Kronecker Factorization Theorem for the exceptional ten dimensional Kac superalgebra $K_{10}$. Also, Y. Popov [15] studied the representations of simple finite-dimensional noncommutative Jordan superalgebras and proved some analogues of the Kronecker factorization theorem for such superalgebras.
The aim of this paper is to determine the structure of Malcev (super)algebras containing the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$. So the paper is organized as follows: In Section 2, we provided definitions about Malcev (super)algebras and their representations. Also we describe some special subvariety $\mathcal{H}$ of Malcev (super)algebras and their representations. In Section 3, we prove analogs of the Kronecker Factorization Theorems for Malcev (super)algebras that contain $\mathbb{M}$. In Section 4, we establish certain equivalences of categories. Finally, in section $5$, we study (super)algebras and modules with (super)involutions.
@@ 2.2 Representation of Malcev superalgebras
given by $\alpha(z)=z\tilde{\alpha}(v,x,y)$. Then $V\simeq\mathbb{M}\alpha$ as modules.
@@ 3 Factorization Theorems
In this section we consider arbitrary Malcev algebras $\mathcal{M}$ containing the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$ such that $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. Also, we consider arbitrary Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ whose even part $\mathcal{M}_{0}$ contains $\mathbb{M}$ with $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. In the general context, we can drop this assumption and so we describe the Malcev (super)algebras containing $\mathbb{M}$.
@@ 3.1 Factorization Theorem for Malcev algebras
The objective of this subsection is to prove an analogy of the Kronecker Factorization Theorem for Malcev algebras that contain $\mathbb{M}$.
@@
Let $\mathcal{M}$ be a Malcev algebra such that $\mathcal{M}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}\mathcal{U}$ for a certain commutative associative algebra $\mathcal{U}$.
Proof: Consider $\mathcal{M}$ as a Malcev $\mathbb{M}$-module. Then by Corollary 2.7, $\mathcal{M}$ is completely reducible, that is, $\mathcal{M}=\sum_{i}\oplus V_{i}$, where $V_{i}$ is an irreducible almost faithful non-Lie Malcev module over $\mathbb{M}$. By Proposition 2.9 we have $0\neq\alpha_{i}(V_{i},\mathbb{M},\mathbb{M})$ because $\mathbb{M}$ is a non-Lie Malcev algebra. Let us take elements $v_{i}\in V_{i}$ and $a_{i},b_{i}\in\mathbb{M}$ such that
@@
where for each $i$ we have $V_{i}\simeq\mathbb{M}\alpha_{i}$. Also by Theorem 2.6, $\mathcal{M}$ is an $\mathcal{H}$-module for $\mathbb{M}$, then $E=\mathbb{M}\oplus\mathcal{M}$ is a Malcev algebra in the variety $\mathcal{H}$; this implies that $\mathcal{M}$ satisfies the identity $h=0$. So, by Lemma 2.2 for all $x,y\in\mathcal{M}$
@@
Let $\mathcal{M}$ be a Malcev algebra containing $\mathbb{M}$. Then $\mathcal{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}A)\oplus\mathcal{N}$, where $A$ is a certain commutative associative algebra.
@@
On the other hand, from Corollary 2.4 $\mathcal{M}=N_{\mathcal{M}}\oplus J_{\mathcal{M}}$, where the submodules $N_{\mathcal{M}}$ and $J_{\mathcal{M}}$ are given by $N_{\mathcal{M}}=\{x\in\mathcal{M}:J(x,\mathbb{M},\mathbb{M})=0\}$ and $J_{\mathcal{M}}=J(\mathcal{M},\mathbb{M},\mathbb{M})$. Moreover, by (ii) of Corollary 2.4 $N_{\mathcal{M}}\mathbb{M}=0$ and $J_{\mathcal{M}}$ is a direct sum of the copies of the adjoint module of $\mathbb{M}$, that is, $J_{\mathcal{M}}=\sum_{j}\oplus\textup{Reg}_{j}\,\mathbb{M}$, where $\textup{Reg}_{j}\,\mathbb{M}\cong\textup{Reg}\,\mathbb{M}$ for all $j$. So $J_{\mathcal{M}}=\sum_{j}\oplus\mathbb{M}\alpha_{j}$, where $\alpha_{j}\in\Gamma(\mathcal{M})$ and they are some of the elements that appear in the first decomposition of $\mathcal{M}$. Denote $A=\sum_{j}F\alpha_{j}$ the span of all $\alpha_{j}$. Then $J_{\mathcal{M}}=\mathbb{M}A.$
@@ 3.2 Factorization Theorem for Malcev superalgebras
The objective of this subsection is to prove an analogy of the Kronecker Factorization Theorem for Malcev superalgebras whose even part contains the 7-dimensional exceptional Malcev algebra $\mathbb{M}$.
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra such that $\mathcal{M}_{0}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}U$ for a certain supercommutative associative superalgebra $U$.
@@
because for each $i$ we have $V_{i}\simeq\mathbb{M}\alpha_{i}$. As in the Theorem 3.1 $\alpha_{i}\in\widetilde{\Gamma}_{0}(\mathcal{M})$.
@@
because for each $j$ we have $W_{j}\simeq\mathbb{M}\beta_{j}$, with
@@
Using Theorem 3.3 we can drop the assumption that $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. The proof of the Theorem 3.4 mimics the Theorem 3.2 proof of the corresponding result for Malcev algebras that contains the $7$-dimensional exceptional Malcev algebra.
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra such that $\mathcal{M}_{0}$ contains $\mathbb{M}$. Then $\mathcal{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}\mathcal{A})\oplus\mathcal{N^{\prime}}$, where $\mathcal{A}$ is a certain supercommutative associative superalgebra.
@@ 4 Some equivalence of categories
In this section, we provide some equivalence of categories. In the case of Malcev algebras, observe that when $\mathbb{M}=\mathcal{M}_{7}(F)$, the Theorem 3.1 states that the Malcev algebra $\mathcal{M}$ is coordinated by $\mathcal{U}$, that is, $\mathcal{M}\cong\mathcal{M}_{7}(\mathcal{U})$. Thus the following result establishes that in this case $\mathcal{M}$ is Morita equivalent to $\mathcal{U}$.
@@
  • let $\textbf{\mbox{Mod}}_{\textbf{\mbox{Malcev}}}\,\mathcal{M}_{7}(\mathcal{U})$ denote the category of Malcev modules $V$ over $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$.
  • let $\textbf{\mbox{Mod}}_{\textbf{\mbox{Ass-Com}}}\,\mathcal{U}$ denote the category of unital associative and commutative modules $W$ over $\mathcal{U}$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$.
We define a functor $\mathbf{T}$ from the category $\textbf{\mbox{Mod}}_{\textbf{\mbox{Ass-Com}}}\,\mathcal{U}$ into the category $\textbf{\mbox{Mod}}_{\textbf{\mbox{Malcev}}}\,\mathcal{M}_{7}(\mathcal{U}).$
Let $W\in\mbox{Obj}(\textbf{\mbox{Mod}}_{\textbf{\mbox{Ass-Com}}}\,\mathcal{U})$ and $E=\mathcal{U}+W$ be the split null extension of $\mathcal{U}$ by $W$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$, $E$ is a unital associative and commutative algebra. Consider the Cayley-Dickson matrix algebra $C(E)$ that contains $C(\mathcal{U})$ as a subalgebra. Also $C(E)$ contains the ideal $C(W)$ which is the set of matrices of $C(E)$ whose entries are in the ideal $W$ of $E$. So introducing the commutator in $C(E)$ we obtain the non-Lie Malcev algebra
@@
that contains the non-Lie Malcev algebra $\mathcal{M}_{7}(\mathcal{U})=C(\mathcal{U})^{(-)}/F\cdot e_{0}$. Moreover $\mathcal{M}_{7}(E)$ contains the ideal $V=\mathcal{M}_{7}(W)$. Then $V=\mathcal{M}_{7}(W)$ is a Malcev module over $\mathcal{M}_{7}(\mathcal{U})$ relative to the multiplication defined in $\mathcal{M}_{7}(E)$. Note that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$ because $W$ is unital. We call to $V$ the $\mathcal{M}_{7}(\mathcal{U})$-module associated with the given module $W$ of $\mathcal{U}$ and denote
@@
$\mathbf{T}:\textbf{\mbox{Mod}}_{\textbf{\mbox{Ass-Com}}}\,\mathcal{U}\longrightarrow\textbf{\mbox{Mod}}_{\textbf{\mbox{Malcev}}}\,\mathcal{M}_{7}(\mathcal{U})$
is really a functor. In addition each pair of objects $W$ and $W^{\prime}$ of $\textbf{\mbox{Mod}}_{\textbf{\mbox{Ass-Com}}}\,\mathcal{U}$, the following equality is valid:
@@
To complete our reduction of the theory of modules for $\mathcal{M}_{7}(\mathcal{U})$ to that of modules for $\mathcal{U}$ we shall now show that every $\mathcal{M}_{7}(\mathcal{U})$-module is isomorphic to a module associated with a module for $\mathcal{U}$.
@@
Let $V$ be a Malcev module for $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$. Then there exists a unital commutative and associative module $W$ for $\mathcal{U}$ such that $V$ is isomorphic to $\mathbf{T}(W)$.
Proof: The fact that $V$ is a Malcev module over $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$, implies that in particular $v\mathcal{M}_{7}(\mathcal{U})\neq 0$ for any $0\neq v\in V$. Let $\mathcal{M}=\mathcal{M}_{7}(\mathcal{U})\oplus V$ be the split null extension of $\mathcal{M}_{7}(\mathcal{U})$ by $V$. So $\mathcal{M}$ is a Malcev algebra containing the 7-dimensional simple non-Lie Malcev algebra $\mathcal{M}_{7}(F)$ as a subalgebra such that $m\mathcal{M}_{7}(F)\neq 0$ for any $0\neq m\in\mathcal{M}$, then by Theorem 3.1 there is a certain commutative and associative algebra $D$ such that $\mathcal{M}=\mathcal{M}_{7}(D).$ Let $W$ be the set of the elements of $D$ that appear in the matrix entries of $V$. Thus $V=\mathcal{M}_{7}(W)$ where $W\triangleleft D$ and $W^{2}=0$ in $D$, because $V\triangleleft\mathcal{M}$ and $V^{2}=0$ in $\mathcal{M}$; so $D=\mathcal{U}\oplus W$. Then $D$ is a split null extension of $\mathcal{U}$ by its module $W$, hence $W$ is a commutative and associative module over $\mathcal{U}$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$. Therefore $\mathbf{T}(W)=V.$
@@
The categories $\textbf{{Mod}}_{\textbf{{Ass-Com}}}\,\mathcal{U}$ and $\textbf{{Mod}}_{\textbf{{Malcev}}}\,\mathcal{M}_{7}(\mathcal{U})$ are Morita equivalent.
Now in the case of Malcev superalgebras, the Theorem 3.3 implies that if $\mathbb{M}=\mathcal{M}_{7}(F)$, then the Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ is coordinated by $U=U_{0}\oplus U_{1}$, that is, $\mathcal{M}\cong\mathcal{M}_{7}(U)$. Thus the following result establishes that in this case $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ is Morita equivalent to $U=U_{0}\oplus U_{1}$.
@@
  • let $\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Malcev}}}\,\mathcal{M}_{7}(U)$ denote the category of Malcev modules $V=V_{0}\oplus V_{1}$ over $\mathcal{M}_{7}(U)$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any homogeneous element $0\neq v\in V_{0}\cup V_{1}$.
  • let $\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Ass-Com}}}U$ denote the category of unital associative supercommutative supermodules $W=W_{0}\oplus W_{1}$ over $U$ such that $wU\neq 0$ for any homogeneous element $0\neq w\in W_{0}\cup W_{1}$.
@@
$\mathbf{S}:\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Ass-Com}}}U\longrightarrow\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Malcev}}}\,\mathcal{M}_{7}(U).$
@@
The categories $\widetilde{\textbf{{Mod}}}_{\textbf{{Ass-Com}}}U$ and $\widetilde{\textbf{{Mod}}}_{\textbf{{Malcev}}}\,\mathcal{M}_{7}(U)$ are Morita equivalent.
@@
denote the non-singular symmetric bilinear form defined in $E$. Then consider the associative algebra with involution $\mathcal{L}(E)$ where

### v3

@@ Abstract
We prove that a Malcev algebra $\mathcal{M}$ which belongs to a certain variety of Malcev algebras $\mathcal{H}$ such that $\mathcal{M}$ contains the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}A)\oplus\mathcal{N}$, where $A$ is a commutative associative algebra. Also, we prove that a Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ in $\mathcal{H}$ whose even part $\mathcal{M}_{0}$ contains $\mathbb{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}\mathcal{A})\oplus\mathcal{N^{\prime}}$, where $\mathcal{A}$ is a supercommutative associative superalgebra.
@@ 1 Introduction
In the case of superalgebras, M. López-Díaz and I. Shestakov [14, 13] studied the representations of simple alternative and exceptional Jordan superalgebras in characteristic $3$ and through these representations, they obtained some analogues of the Kronecker Factorization Theorem for these superalgebras. Also, the author [15] obtained some analogues of the Kronecker Factorization Theorem for central simple alternative superalgebras $M_{(1\mid 1)}(F)$, $\mathbb{O}(4|4)$ and $\mathbb{O}[u]$, where in particular the Kronecker Factorization Theorem for $M_{(1\mid 1)}(F)$ answers the analogue for superalgebras of the Jacobson’s problem [10]. We state that Jacobson’s problem [10] was recently solved by the author and I. Shestakov [16] in the split case. Similarly, C. Martinez and E. Zelmanov [17] obtained a Kronecker Factorization Theorem for the exceptional ten dimensional Kac superalgebra $K_{10}$. Also, A. Pozhidaev and I. Shestakov [20] studied the representations of simple finite-dimensional noncommutative Jordan superalgebras and proved some analogues of the Kronecker factorization theorem for such superalgebras (see also [19]).
In [5, (6)], Filippov introduced a very interesting identity and considered (see [6, 7, 8, 9]) the subvariety $\mathcal{H}$ of those Malcev algebras satisfying this identity. He proved in [7] that, in a sense, the subvariety $\mathcal{H}$ plays the same role as the variety of Lie algebras in the structure theory of Malcev algebras. Moreover, the central closures of the prime algebras in $\mathcal{H}$ and the simple algebras in $\mathcal{H}$ are the 3-dimensional simple Lie algebra $\mathfrak{sl}(2,F)$ and the 7-dimensional exceptional Malcev algebra $\mathbb{M}$ (see [3]).
The aim of this paper is to determine the structure of Malcev (super)algebras in $\mathcal{H}$ containing the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$. So the paper is organized as follows: In Section 2, we provided definitions about Malcev (super)algebras and their representations. Also we describe the special subvariety $\mathcal{H}$ of Malcev (super)algebras and their representations. In Section 3, we prove some Kronecker Factorization Theorems for Malcev (super)algebras in $\mathcal{H}$ that contain $\mathbb{M}$. In Section 4, we establish certain equivalences of categories. Finally, in section $5$, we study (super)algebras and modules with (super)involutions.
@@ 2.2 Representation of Malcev superalgebras
given by $\alpha(z)=z\tilde{\alpha}(v,x,y)$. Then $V\cong\mathbb{M}\alpha$ as modules.
@@ 3 Factorization Theorems
In this section we consider Malcev algebras $\mathcal{M}$ in $\mathcal{H}$ containing the $7$-dimensional exceptional Malcev algebra $\mathbb{M}$ such that $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. Also, we consider arbitrary Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ in the variety $\mathcal{H}$ whose even part $\mathcal{M}_{0}$ contains $\mathbb{M}$ with $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. In the general context, we can drop this assumption and so we describe the Malcev (super)algebras containing $\mathbb{M}$.
@@ 3.1 Factorization Theorem for Malcev algebras
The objective of this subsection is to prove an analogue of the Kronecker Factorization Theorem for Malcev algebras in $\mathcal{H}$ that contain $\mathbb{M}$.
@@
Let $\mathcal{M}$ be a Malcev algebra in $\mathcal{H}$ such that $\mathcal{M}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}\mathcal{U}$ for a certain commutative associative algebra $\mathcal{U}$.
Proof: Consider $\mathcal{M}$ as a Malcev $\mathbb{M}$-module. Then by Corollary 2.7, $\mathcal{M}$ is completely reducible, that is, $\mathcal{M}=\sum_{i}\oplus V_{i}$, where $V_{i}$ is an irreducible almost faithful non-Lie Malcev module over $\mathbb{M}$, that is, $V_{i}\cong\textup{Reg}_{i}\,\mathbb{M}\cong\textup{Reg}\,\mathbb{M}$ for all $i$. Also by Proposition 2.9 we have $0\neq\alpha_{i}(V_{i},\mathbb{M},\mathbb{M})$ because $\mathbb{M}$ is a non-Lie Malcev algebra. Let us take elements $v_{i}\in V_{i}$ and $a_{i},b_{i}\in\mathbb{M}$ such that
@@
where for each $i$ we have $\textup{Reg}_{i}\,\mathbb{M}\cong\mathbb{M}\alpha_{i}$. Moreover, since $\mathcal{M}$ satisfies $H(\mathcal{M})=0,$ then by Lemma 2.2 for all $x,y\in\mathcal{M}$ we have
@@
Let $\mathcal{M}$ be a Malcev algebra in $\mathcal{H}$ containing $\mathbb{M}$. Then $\mathcal{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}A)\oplus\mathcal{N}$, where $A$ is a certain commutative associative algebra.
@@
On the other hand, from Corollary 2.4, $\mathcal{M}=N_{\mathcal{M}}\oplus J_{\mathcal{M}}$, where the submodules $N_{\mathcal{M}}$ and $J_{\mathcal{M}}$ are given by $N_{\mathcal{M}}=\{x\in\mathcal{M}:J(x,\mathbb{M},\mathbb{M})=0\}$ and $J_{\mathcal{M}}=J(\mathcal{M},\mathbb{M},\mathbb{M})$. Moreover, by (ii) of Corollary 2.4, $N_{\mathcal{M}}\mathbb{M}=0$ and $J_{\mathcal{M}}$ is a direct sum of the copies of the adjoint module of $\mathbb{M}$, that is, $J_{\mathcal{M}}=\sum_{j}\oplus\textup{Reg}_{j}\,\mathbb{M}$, where $\textup{Reg}_{j}\,\mathbb{M}\cong\textup{Reg}\,\mathbb{M}$ for all $j$. So $J_{\mathcal{M}}=\sum_{j}\oplus\mathbb{M}\alpha_{j}$, where $\alpha_{j}\in\Gamma(\mathcal{M})$ and they are some of the elements that appear in the first decomposition of $\mathcal{M}$. Denote $A=\sum_{j}F\alpha_{j}$ the span of all $\alpha_{j}$. Then $J_{\mathcal{M}}=\mathbb{M}A.$
@@ 3.2 Factorization Theorem for Malcev superalgebras
The objective of this subsection is to prove an analogue of the Kronecker Factorization Theorem for Malcev superalgebras in $\mathcal{H}$ whose even part contains the 7-dimensional exceptional Malcev algebra $\mathbb{M}$.
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra in $\mathcal{H}$ such that $\mathcal{M}_{0}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}U$ for a certain supercommutative associative superalgebra $U$.
@@
because for each $i$ we have $V_{i}\cong\mathbb{M}\alpha_{i}$. As in the Theorem 3.1, $\alpha_{i}\in\widetilde{\Gamma}_{0}(\mathcal{M})$.
@@
because for each $j$ we have $W_{j}\cong\mathbb{M}\beta_{j}$, with
@@
Using Theorem 3.3 we can drop the assumption that $m\mathbb{M}\neq 0$ for any homogeneous element $0\neq m\in\mathcal{M}_{0}\cup\mathcal{M}_{1}$. The proof of the Theorem 3.4 mimics the Theorem 3.2 proof of the corresponding result for Malcev algebras in $\mathcal{H}$ that contains the $7$-dimensional exceptional Malcev algebra.
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra in $\mathcal{H}$ such that $\mathcal{M}_{0}$ contains $\mathbb{M}$. Then $\mathcal{M}$ is isomorphic to a direct sum $(\mathbb{M}\otimes_{F}\mathcal{A})\oplus\mathcal{N^{\prime}}$, where $\mathcal{A}$ is a certain supercommutative associative superalgebra.
@@ 4 Some equivalence of categories
In this section, we provide some equivalence of categories. In the case of Malcev algebras, observe that when $\mathbb{M}=\mathcal{M}_{7}(F)$, the Theorem 3.1 states that the Malcev algebra $\mathcal{M}$ in $\mathcal{H}$ is coordinated by $\mathcal{U}$, that is, $\mathcal{M}\cong\mathcal{M}_{7}(\mathcal{U})$. Thus the following result establishes that in this case $\mathcal{M}$ is Morita equivalent to $\mathcal{U}$.
@@
  • let $\mathcal{H}-\textbf{\mbox{Mod}}_{\textbf{\mbox{Malc}}}\,\mathcal{M}_{7}(\mathcal{U})$ denote the category of $\mathcal{H}$-modules $V$ over $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$.
  • let $\textbf{\mbox{Mod}}_{\textbf{\mbox{Assoc-Com}}}\,\mathcal{U}$ denote the category of unital associative and commutative modules $W$ over $\mathcal{U}$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$.
We define a functor $\mathbf{T}$ from the category $\textbf{\mbox{Mod}}_{\textbf{\mbox{Assoc-Com}}}\,\mathcal{U}$ into the category $\mathcal{H}-\textbf{\mbox{Mod}}_{\textbf{\mbox{Malc}}}\,\mathcal{M}_{7}(\mathcal{U}).$
Let $W\in\mbox{Obj}(\textbf{\mbox{Mod}}_{\textbf{\mbox{Assoc-Com}}}\,\mathcal{U})$ and $E=\mathcal{U}+W$ be the split null extension of $\mathcal{U}$ by $W$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$, $E$ is a unital associative and commutative algebra. Consider the Cayley-Dickson matrix algebra $C(E)$ that contains $C(\mathcal{U})$ as a subalgebra. Also $C(E)$ contains the ideal $C(W)$ which is the set of matrices of $C(E)$ whose entries are in the ideal $W$ of $E$. So introducing the commutator in $C(E)$ we obtain the non-Lie Malcev algebra
@@
that contains the non-Lie Malcev algebra $\mathcal{M}_{7}(\mathcal{U})=C(\mathcal{U})^{(-)}/F\cdot e_{0}$. Moreover, since $\mathcal{M}_{7}(F)$ belongs to $\mathcal{H}$ and $\mathcal{M}_{7}(E)=\mathcal{M}_{7}(F)\otimes_{F}E,$ then $\mathcal{M}_{7}(E)$ belongs to $\mathcal{H}$ and contains the ideal $V=\mathcal{M}_{7}(W)$. Therefore, $V=\mathcal{M}_{7}(W)$ is a $\mathcal{H}$-module over $\mathcal{M}_{7}(\mathcal{U})$ relative to the multiplication defined in $\mathcal{M}_{7}(E)$. Note that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$ because $W$ is unital. We call to $V$ the $\mathcal{M}_{7}(\mathcal{U})$-module associated with the given module $W$ of $\mathcal{U}$ and denote
@@
$\mathbf{T}:\textbf{\mbox{Mod}}_{\textbf{\mbox{Assoc-Com}}}\,\mathcal{U}\longrightarrow\mathcal{H}-\textbf{\mbox{Mod}}_{\textbf{\mbox{Malc}}}\,\mathcal{M}_{7}(\mathcal{U})$
is really a functor. In addition each pair of objects $W$ and $W^{\prime}$ of $\textbf{\mbox{Mod}}_{\textbf{\mbox{Assoc-Com}}}\,\mathcal{U}$, the following equality is valid:
@@
To complete our reduction of the theory of $\mathcal{H}$-modules for $\mathcal{M}_{7}(\mathcal{U})$ to that of modules for $\mathcal{U}$ we shall now show that every $\mathcal{H}$-module over $\mathcal{M}_{7}(\mathcal{U})$ is isomorphic to a module associated with a module for $\mathcal{U}$.
@@
Let $V$ be a $\mathcal{H}$-module for $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$. Then there exists a unital associative and commutative module $W$ for $\mathcal{U}$ such that $V$ is isomorphic to $\mathbf{T}(W)$.
Proof: The fact that $V$ is a $\mathcal{H}$-module over $\mathcal{M}_{7}(\mathcal{U})$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any $0\neq v\in V$, implies that in particular $v\mathcal{M}_{7}(\mathcal{U})\neq 0$ for any $0\neq v\in V$. Let $\mathcal{M}=\mathcal{M}_{7}(\mathcal{U})\oplus V$ be the split null extension of $\mathcal{M}_{7}(\mathcal{U})$ by $V$. So $\mathcal{M}$ is a Malcev algebra in $\mathcal{H}$ containing the 7-dimensional simple non-Lie Malcev algebra $\mathcal{M}_{7}(F)$ as a subalgebra such that $m\mathcal{M}_{7}(F)\neq 0$ for any $0\neq m\in\mathcal{M}$, then by Theorem 3.1 there is a certain commutative and associative algebra $D$ such that $\mathcal{M}=\mathcal{M}_{7}(D).$ Let $W$ be the set of the elements of $D$ that appear in the matrix entries of $V$. Thus $V=\mathcal{M}_{7}(W)$ where $W\triangleleft D$ and $W^{2}=0$ in $D$, because $V\triangleleft\mathcal{M}$ and $V^{2}=0$ in $\mathcal{M}$; so $D=\mathcal{U}\oplus W$. Then $D$ is a split null extension of $\mathcal{U}$ by its module $W$, hence $W$ is an associative and commutative module over $\mathcal{U}$ such that $w\mathcal{U}\neq 0$ for any $0\neq w\in W$. Therefore $\mathbf{T}(W)=V.$
@@
The categories $\textbf{{Mod}}_{\textbf{{Assoc-Com}}}\,\mathcal{U}$ and $\mathcal{H}-\textbf{{Mod}}_{\textbf{{Malc}}}\,\mathcal{M}_{7}(\mathcal{U})$ are Morita equivalent.
Now in the case of Malcev superalgebras, the Theorem 3.3 implies that if $\mathbb{M}=\mathcal{M}_{7}(F)$, then the Malcev superalgebra $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ in $\mathcal{H}$ is coordinated by $U=U_{0}\oplus U_{1}$, that is, $\mathcal{M}\cong\mathcal{M}_{7}(U)$. Thus the following result establishes that in this case $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ is Morita equivalent to $U=U_{0}\oplus U_{1}$.
@@
  • let $\mathcal{H}-\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Malc}}}\,\mathcal{M}_{7}(U)$ denote the category of $\mathcal{H}$-modules $V=V_{0}\oplus V_{1}$ over $\mathcal{M}_{7}(U)$ such that $v\mathcal{M}_{7}(F)\neq 0$ for any homogeneous element $0\neq v\in V_{0}\cup V_{1}$.
  • let $\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Assoc-Com}}}U$ denote the category of unital associative supercommutative supermodules $W=W_{0}\oplus W_{1}$ over $U$ such that $wU\neq 0$ for any homogeneous element $0\neq w\in W_{0}\cup W_{1}$.
@@
$\mathbf{S}:\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Assoc-Com}}}U\longrightarrow\mathcal{H}-\widetilde{\textbf{\mbox{Mod}}}_{\textbf{\mbox{Malc}}}\,\mathcal{M}_{7}(U).$
@@
The categories $\widetilde{\textbf{{Mod}}}_{\textbf{{Assoc-Com}}}U$ and $\mathcal{H-}\widetilde{\textbf{{Mod}}}_{\textbf{{Malc}}}\,\mathcal{M}_{7}(U)$ are Morita equivalent.
@@
denote the non-singular symmetric bilinear form defined in $E$. Then consider the associative algebra with involution $\mathcal{L}(E)$, where